40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 16 10:15:02 2009, steve, Update, IOO, IP ang & pos recentered ip4d.jpg
    Reply  Thu Oct 22 08:45:58 2009, steve, Update, IOO, IP ang & pos recentered pointing1000.jpg
Message ID: 2132     Entry time: Thu Oct 22 08:45:58 2009     In reply to: 2102
Author: steve 
Type: Update 
Category: IOO 
Subject: IP ang & pos recentered 

Quote:

Pointing stability of 4 days. Initial pointing does not go through suspended optics. It is launched  right after the Piezo Jena steering mirrors in the BS chamber.

IP-ANG on epics screen is  C1:ASC-IBQPD_X and Y in dataviewer  were recentered. This beam is clipping a bit in ETMX chamber  pick off mirror.

IP-POS pick  off is in the BS chamber and it's qpd on the BS_ISCT This beam is also clipping just a little bit. This is easy to fix. We'll have to remove an iris from the BS optical levers table.

note: arms were not locked when I recentered

 IP-ANG clipping can be traced back to our last vent of Aug. 18, 2008  See elog entry #845

This was an after earth quake - sus repair vent

Attachment 1: pointing1000.jpg  123 kB  | Show | Show all
ELOG V3.1.3-